OpenStudy (anonymous):

6(a + 3) = 21 – 3(1 – 2a) our answers could be no solution, all real numbers, –3, –1,

OpenStudy (anonymous):

6(a+3)=21-3(1-2a) 6a+18=3-6a 12a+18=3 12a=-15 a=-.8 so I'm guessing all real numbers, it's the only one that makes sense?